OpenStudy (anonymous):

Write the equation of the function g(x) if g(x) = f(x – 2) +4 and f(x) = x^3 + 2.

jimthompson5910 (jim_thompson5910):

f(x) = x^3 + 2 f(x) = (x)^3 + 2 f(x-2) = (x-2)^3 + 2 ... replace every x with x-2 f(x-2) = ???

OpenStudy (anonymous):

Would you add the 4 along with the (x-2)^3 + 2? f(x-2) = (x-2)^3 + 6? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it should be f(x-2)+4 = (x-2)^3 + 6, GRoa1, but you have the right idea

jimthompson5910 (jim_thompson5910):

optionally you expand out (x-2)^3 and then combine like terms to simplify

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Mmmm, I am confused now. How do you get g(x) from f(x-2) = X^3 - 6x^2 + 12x -2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

because at the top it says that g(x) = f(x – 2) +4